OpenStudy (anonymous):

How much do you need to deposit today if you want to have one million dollars in 20 years at an account which pays 6% interest compounded monthly?

OpenStudy (kropot72):

Let the amount of the required deposit be P. Then the value of P can be found as follows: \[\large P=\frac{10^{6}}{1.005^{(20\times12)}}=you\ can\ calculate\]

OpenStudy (anonymous):

how did you get 1.005?

OpenStudy (kropot72):

0.005 is the effective monthly rate of interest as a decimal. It is the value of 0.06/12. Then 0.005 must be added to 1, the reason being to follow the equation for monthly compounding: \[\large A=P(1+\frac{r}{12})^{(n \times12)}\] where r is the annual rate as a decimal, and n is the number of years.

OpenStudy (anonymous):

makes sense now! thank you! can you help me with another one?

OpenStudy (kropot72):

Please post any more as new questions. Thank you :)